given:5+5(x+6)=50
prove:5x+9=24
make a two column proof

Answer by jim_thompson5910(35256)

StatementReason
1. 5+5(x+6)=50 Given
2. 5+5x+30=50 Distributive Property
3. 5x+35=50 Combining Like Terms
4. 5x+9+26=50 Breaking up 35 into 9+26
5. 5x+9+26+(-26)=50+(-26) Additive Property of Equality
6. 5x+9+(26+(-26))=50+(-26) Associative Property of Addition
7. 5x+9+0=24 Adding terms
8. 5x+9=24 Additive Identity Property